13: 9781285868806

pactice Exercises EE61 p 235 PE6-1A Flexible budgeting At the beginning of the period, the Assembly Department budgeted direct labor and property tax of $12,000 for 7,000 hours of production. The department pleted 7,500 hours of production. Determine the budget for the department, ass that it uses flexible budgeting. OBJ, 2 of $112,000 actually com- uming EE61 p25 PE6-1B Flexible budgeting At the beginning of the period, the Fabricating Department budgeted $9,280 and equipment depreciation of $2,300 for 640 hours of production. actually completed 600 hours of production. Determine the budget for the de OBJ. 2 direct labor of The department assuming that it uses flexible budgeting. EE 6-2 p 238 PE 6-2A MyLife Chronicles Inc. projected sales of 240,000 diaries for 2016. The estimated January 1 2016, inventory is 19,900 units, and the desired December 31, 2016, inventory is 18,800 units. What is the budgeted production (in units) for 2016 Production budget OBJ. 4 EE 6-2 p 238 PE 6-2B Production budget OBJ. 4 Magnolia Candle Inc. projected sales of 75,000 candles for 2016. The estimated January 1 2016, inventory is 3,500 units, and the desired December 31, 2016, inventory is 2,700 units, What is the budgeted production (in units) for 2016? EE6-3 p 240 PE 6-3A Direct materials purchases budget OBJ.4 MyLife Chronicles Inc. budgeted production of 238,900 diaries in 2016. Paper is required to produce a diary. Assume five square yards of paper are required for each diary. The estimated January 1, 2016, paper inventory is 32,400 square yards. The desired December 31, 2016, paper inventory is 30,800 square yards. If paper costs $0.30 per square yard, determine the direct materials purchases budget for 2016. PE 6-3BDirect materials purchases budget Magnolia Candle Inc. budgeted production of 74,200 candles in 2016. Wax is required to produce a candle. Assume eight ounces (one-half of a pound) of wax is required for each candle. The estimated January 1, 2016, wax inventory is 2,500 pounds. The desired December 31, 2016, wax inventory is 2,100 pounds. If candle wax costs $4.10 per pound, determine the direct materials purchases budget for 2016. EE 6-3 p 240 ORI Now
Answer #1

Answers

* From next time, please mention which problem you want specifically.

  • PE 6-1A

Budget for Department = ($112000/7000 hours) x 7500 hours + $ 12000
= 120000 + 12000
= $ 132,000 = Answer

  • PE 6-1B

Budget for Department = ($9280/640 hours) x 600 hours + $ 2300
= 8700 + 2300
= $ 11,000 = Answer

  • PE 6-2A

Budgeted production = Estimated sale + Desired ending inventory – Beginning inventory.
= 240000 + 18800 – 19900
= 238,900

  • PE 6-2B
    Budgeted production = Estimated sale + Desired ending inventory – Beginning inventory.
    = 75000 + 2700 – 3500
    = 74,200 units
  • PE 6-3A

A

Budgeted production

               238,900

B

Square yards required per unit

                            5

C = A x B

Total requirements

           1,194,500

D

Desired ending inventory

                 30,800

E

Beginning Inventory

                 32,400

F = C+D-E

Total square yards to be purchased

           1,192,900

G

Cost per Square yards

$                  0.30

H = F x G

Budgeted direct material purchase cost

$           357,870

  • PE 6-3B

A

Budgeted production

                 74,200

B

Wax required per unit (pounds)

                      0.50

C = A x B

Total requirements

                 37,100

D

Desired ending inventory

                   2,100

E

Beginning Inventory

                   2,500

F = C+D-E

Total wax to be purchased

                 36,700

G

Cost per pounds

$                  4.10

H = F x G

Budgeted direct material purchase cost

$           150,470